Find the slope and the -y intercept of the line
8x + 3y = 41

Answer by lwsshak3(11628)

Find the slope and the -y intercept of the line
8x + 3y = 41
equation of a line: y=mx+b, m=slope, b=y-intercept
For given equation:
8x+3y=41
3y=-8x+41
y=-8x/3+41/3
slope,m=-8/3
y-intercept=(0,41/3)